OpenStudy (compassionate):

34,100 at 4% for 3 years.

OpenStudy (compassionate):

@Miracrown @Zale101 I would say 3(34,100 x 0.4) Right?

OpenStudy (tainted):

34100 x 0.04

OpenStudy (tainted):

Because it isn't 10 or over it would be 0.04 instead of 0.4(40%)

OpenStudy (compassionate):

So if we do 34100 x 0.04 we get: 4092 Now 341,000 + 4092 = 38.192

OpenStudy (tainted):

Correct, but you do mean 34,100 instead of 341,000 right?

OpenStudy (compassionate):

Yes. Hehe.

OpenStudy (compassionate):

4,000 at 3% for 4 years 4,000 x 0.03 * 4 120 * 40 = 480 4,000 + 480 = 4,480

OpenStudy (tainted):

3(34,100 * 4%) (34,100 x 0.04) = 1364 1364 * 3 = 4092 34,100+4092= 38,192

OpenStudy (tainted):

Wait, you changed around the 4% and the 3 years from your original topic question.

OpenStudy (compassionate):

IT's fine! :) This is a new one

OpenStudy (tainted):

Oh, well yes then you would do the same thing, just reversed.

OpenStudy (compassionate):

19,00 at 5.9% for 2(3/4) years

Miracrown (miracrown):

The principal amount is 19000, is that 19,000 or 1900?

Miracrown (miracrown):

Can you confirmm???

OpenStudy (compassionate):

1,900 my computer is being stupid sorry :(

Miracrown (miracrown):

So? -.- My laptop has heating problems.

OpenStudy (compassionate):

Mine has typing problems.

Miracrown (miracrown):

Mine has problems with EVERYTHING! >.<

OpenStudy (compassionate):

Could we say: 19,00 at 5.9% for 2(3/4) years Convert 3/4 into a decimal, multiply it by 2, and do it that way?

Miracrown (miracrown):